OpenStudy (ellaella02):

NEED HELP PLEASE What is the simplified version of sqrt p^7

OpenStudy (mww):

We use two index laws and combine them, the power of a power rule and the sqrt as a fractional index of half \[(a^m)^n = a^{mn}\] \[\sqrt a = a^{\frac{ 1 }{ 2 }}\] \[\sqrt{ a^7} = (a^7)^{\frac{ 1 }{ 2 }} = a^{ \frac{ 7 }{ 2 }}\] General rule (fractional indices) \[a^\frac{ m }{ n } = (a^m)^\frac{ 1 }{ n } = \sqrt[n]{a^m} ~OR~ (\sqrt[n]{a})^m\]

OpenStudy (kaithang):

Do you still need help?

OpenStudy (mrs.ambrose614):

Did I help you????